AGW1 Alignment with LUCI@ARGOS 20151023 UT

Observers: JHill, DMiller (Tucson)
Telescope Operator: GBechetti (LBTO)
AO Support: DMiller
Telescope Support: JLittle (LBTO)
Software Support: TSargent, DCox
Instrument Support: ESolheid (LBTO)
Engineering Support: DAshby

Summary

Fog, then Az axis problems, then Laser alignment, so didn't get on a star until 8UT. ......before we started AGW1 alignment.

Had some troubles fighting new features in GCS - it knows when we adjust the star off the old hotspot. We got the new guiding hotspot precisely adjusted (within 0.5 pixel, limited by dome seeing). We got the probe roughly adjusted to the rotator center. There is still something loose in the AGW as the circles have steps of about 2 arcsec as we rotate. We got a rough adjustment of the tertiary tip-tilt (reducing the wobble to ~2/3 subap from several subaps). The problem there was that we ran out of negative tip travel. This probably needs some more negative tilt to gain a bit of tip.

Preparations

John had forgotten to update the PSF/SXSMLUCIFERACollimation.dat file, so did that on the fly.
Added in the global offsets X=+0.216, Y=-0.411 Z=0 RX=+10.8 RY=+5.3
(later updated PSF/SXSMIRTCACollimation.dat)

08:05 First ACQUIRE rpreset after unlocking rotator.

Found star initially at IE=-39 CA=+38. Initial focus was off by -3 mm - not sure why.

08:40 First Active preset. Seeing is about 0.6 arcsec.

AGW1 Guiding Hotspot Alignment

New feature of GCS stops/pauses wfs when you adjust hotspot. That really kills this adjustment, so we had to adjust two params in LUCI_L.cfg to make blue circle bigger. maxOffsetToHotspot maxHotspotFromProbe (changed from 0.54 arcsec to 2.54 arcsec). We aren't completely clear about whether we needed to change both of these. Have to stop and restart GCSL to reread this configuration file each time.

Jonathan's lab guiding hotspot was 263,256 pixels. We started with a brighter star, but moved to BS9107 on-axis for the final adjustment.

+X change in guide hotspot moves wfs spots right.
+Y change in guide hotspot moves wfs spots up (maybe).

We adopted 241.0,248.5 pixels for guidecam_hotspot_x, guidecam_hotspot_y by using the mean offset (average_x and average_y in units of fractional subaps) of the Shack spots in Doug's IDL analysis program gcs_wfsc_display.

AGW1 Probe Alignment to Rotator Center

These images from GCSL are written in /lbt/data/share/tcs/telescope.20151023

10:xx guide 0002 x=101.5 y=112 BUT forgot to move probe to 0,612.5 so the measurement is biased. The circle is oval by a couple arcsec, so something is still loose in AGW1. See Issue 5733 (created, but not much information yet).

10:13 guide 0004 x=120.75 y=113 --> 256.5, 265 Need -15 pix X (+24 to ROT homeoffset; 35554 changes to 35578 counts)
Need -16 pix Y (+450 RAD ooffset; 417375 changes to 417825 microns)

logged in as oac, the configuration file is /lbt/oacontrol/current/etc/oacontrol.conf

10:27 guide 0005 x=116.25 y=107.5 --> 247.5 254 (adopted as close enough given the wobbles)

Adjusting M3 to minimize pupil wobble

11:xx starting at SX M3 absolute Tip=-616 Tilt=+291

11:16 M3 absolute Tip=-1315 Tilt=+291

11:42 M3 absolute Tip=-1455 Tilt=+90 (this more negative Tilt is needed to get Tip as far as it went). These numbers were saved in OSS/positions.conf

12:00 Pointing is IE=-49 CA=-33 at this new tip-tilt position.

New on-axis pupil position was installed in LUCI_L.cfg WFScam_hotspot_x 383.0 WFScam_hotspot_y 220.0

12:02 back to ARGOS laser operations.

-- %USERSIG{JohnHill - 2015-10-23}%
Topic revision: r2 - 25 Oct 2015, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback